Suppose you are given the following data on demand for a product. The price

elasticity of demand when price decreases from $9 to $7 is:

Price ($) Quantity Demanded

10            30

9              40

8              50

7              60

6              70


1
Expert's answer
2022-03-01T14:50:36-0500

"PED= \frac{\Delta Q}{\Delta P}\times \frac{P}{Q}"

"PED= \frac{60-40}{7-9}\times \frac{7}{60}"

= "-10\times \frac {7}{60}= -1.2"